OpenStudy (anonymous):

can anyone help me with this problem? 6^2x+6^x-90=0 im confused on how to start it, would you use the quadratic equation???

OpenStudy (immanuelv):

gotta use logs :D

OpenStudy (perl):

(6^x)^2 + (6^x) - 90 = 0

OpenStudy (perl):

this has the form u^2 + u - 90 = 0

OpenStudy (perl):

( u - 9 ) ( u + 10)

OpenStudy (perl):

now plug in , I substituted 6^x for u

OpenStudy (perl):

(6^x - 9 ) (6^x + 10) =0

OpenStudy (perl):

now use zero product property

OpenStudy (perl):

6^x - 9 = 0 or 6^x + 10 = 0

OpenStudy (anonymous):

oh okay thank you! @perl